Exercise 6-6 Variable and Absorption Costing Unit Product Costs and Income Statements [LO6-1, LO6-2] Lynch Company manufactures and sells a single product. The following costs were incurred during the company’s first year of operations: Variable costs per unit: Manufacturing: Direct materials $ 14 Direct labor $ 5 Variable manufacturing overhead $ 1 Variable selling and administrative $ 1 Fixed costs per year: Fixed manufacturing overhead $ 264,000 Fixed selling and administrative $ 174,000 During the year, the company produced 33,000 units and sold 15,000 units. The selling price of the company’s product is $52 per unit.

Required:

1. Assume that the company uses absorption costing: a. Compute the unit product cost. b. Prepare an income statement for the year.

2. Assume that the company uses variable costing: a. Compute the unit product cost. b. Prepare an income statement for the year.
Answer #1
1a
Direct materials 14
Direct labor 5
Variable manufacturing overhead 1
Fixed manufacturing overhead 8 =264000/33000
Unit product cost 28
b
Absorption Costing Income Statement
Sales 780000 =15000*52
Cost of goods sold 420000 =15000*28
Gross margin 360000
Selling and administrative expenses 189000 =174000+(15000*1)
Net operating income (loss) 171000
2a
Direct materials 14
Direct labor 5
Variable manufacturing overhead 1
Unit product cost 20
b
Variable Costing Income Statement
Sales 780000 =15000*52
Variable expenses:
Variable Cost of goods sold 300000 =15000*20
Variable Selling and administrative expenses 15000 =15000*1
315000
Contribution margin 465000
Fixed expenses:
Fixed manufacturing overhead 264000
Fixed Selling and administrative expenses 174000
438000
Net operating income (loss) 27000